# SemHiTok: A Unified Image Tokenizer via Semantic-Guided Hierarchical Codebook for Multimodal Understanding and Generation

### v4

@@ Abstract
In this paper, we introduce SemHiTok, a unified image Tokenizer via Semantic-Guided Hierarchical codebook that provides consistent discrete representations for multimodal understanding and generation. Recently, unified image tokenizers have sparked exploration within research community, which is designed to capture high-level semantic features for understanding and retaining low-level pixel features for generation. Previous works attempt to train a unified image tokenizer by combining loss for semantic distillation and pixel reconstruction. However, due to the differing levels of features prioritized by multimodal understanding and generation, joint training methods face significant challenges in achieving a good trade-off. SemHiTok addresses this challenge through a novel semantic-guided hierarchical codebook, which builds pixel sub-codebooks on a pretrained semantic codebook. This design decouples semantic and pixel both in terms of structure and training strategy, enabling the tokenizer to capture pixel features while retaining its ability to comprehend high-level semantic information. Our experiments demonstrate that SemHiTok achieves SOTA performance in image reconstruction and multimodal understanding under LLaVA-v1.5 setting. Further, we develop a unified MLLM with SemHiTok, which exhibits superior performance across multimodal understanding and generation tasks. For understanding, SemHiTok achieves impressive performance on most benchmarks. For generation, our model achieves SOTA performance on MJHQ30K in unified MLLMs.
@@ 1 Introduction
Figure 1: Illustration of other tokenizers and SemHiTok.
Table 1: Characteristics comparison among existing unified image tokenizer. Previous works attempt to achieve a better trade-off by structural separation or more powerful codebook type. However, these methods do not achieve separate training. Additionally, their specialized codebook structures lack consideration of compatibility with MLLMs. In comparison, our approach offers a better trade-off and is widely compatible with mainstream MLLMs following the next-token prediction paradigm.
Tokenizer VILA-U[54] TokenFlow[36] SDE[57] SemHiTok(Ours)
Codebook Structure RQ Share-Mapping & MSVQ Vanilla Hierarchical
Separate Training
$\times$
$\times$
$\times$
$\checkmark$
Structural Separation
$\times$
$\checkmark$
$\times$
$\checkmark$
Unifid Representation
$\checkmark$
$\times$
$\checkmark$
$\checkmark$
Unified MLLM head
$\times$
$\checkmark$
$\checkmark$
$\checkmark$
Figure 2: Demonstrating the limitations of dual-tokenizer wshen applied to a unified large model.
@@
In recent years, autoregressive models have achieved great success in natural language processing, and have been extended to the multimodal understanding domain, demonstrating immense potential. This triggers researchers’ interest in unified multimodal understanding and generation by employing a single autoregressive framework. To achieve a unified multimodal large model, the key challenge is designing a tokenizer suitable for both multimodal generation and understanding tasks.
@@
However, there is a vast gap in the visual information required for these two tasks. For instance, models from the CLIP[37; 43; 66] family, commonly used in multimodal understanding tasks, tend to lose visual pixel information. On the contrary, the VQGAN[61; 68] family models, often used in autoregressive generation tasks, lack the ability to extract semantic feature for multimodal understanding tasks. This leads to poor performance when a single tokenizer is applied to a unified MLLM[53; 18; 22]. In light of the aforementioned issues, some recent work [54; 36; 57] as shown in Fig. 1, have attempted to incorporate a semantic learning branch into the original VQGAN training pipeline, aiming to obtain a unified tokenizer via joint optimization. However, due to the hybrid architecture (encoder or codebook) and joint training, such approaches require massive amounts of data to achieve the trade-off between semantic features and pixel features. Additionally, VILA-U employs residual quantization(RQ)[19] to improve the capacity of the discrete space, which prohibits unified modeling with a shared head layer for both text and visual. Although, as shown in Fig. 1(b), SDE[57]employs dual specialized encoder, semantic and pixel information still share a single codebook, which leads to the same problem. TokenFlow[36] uses shared mapping to decouple the semantic branch and pixel branch while maintaining the consistency of the codebook index, but joint training still affects the final performance. Janus[52] introduce a dual-encoder method that separate encoders for understanding and generation tasks to address this conflict, but this increases the complexity of handling mixed tasks and does not fundamentally resolve the feature conflict challenge.
@@
A straightforward approach is to use CLIP and VQGAN to extract semantic and pixel information respectively, and the concatenation of thees two tokens sequence is then used as a unified representation. However, this leads to a doubling of the token sequence count or multiplicative expansion in vocabulary size, depending on whether the concatenation is applied along the length or dimension, as shown in Fig. 2. These limitations underscore a fundamental challenge in the field: how to balance semantic-level and pixel-level information effectively, without compromising the ease of integration into MLLM frameworks?
@@
To address this challenge, we propose SemHiTok, a unified image tokenizer that provides consistent feature representations for multimodal understanding and generation tasks through a unique hierarchical codebook design. Inspired by the observation that image patches with the same semantic code tend to have similar pixel feature, we introduce a novelty hierarchical codebook which uses sub-codebook to model the pixel-level space associated with each semantic code, named Semantic-Guided Hierarchical Codebook(SGHC). Unlike existing approaches, SemHiTok support stage-wise training paradigm where each stage exclusively optimizes specific hierarchy level features, allows us to achieve a better trade-off between semantic and pixel feature extraction. In addition, SemHiTok can be seamlessly integrated into existing MLLMs following the next-token paradigm through a simple codebook flattening operation.
@@
Our contributions can be summarized as follows: (1):A novel unified tokenizer that achieves a trade-off between semantic and pixel information, demonstrating outstanding performance in both image reconstruction and multimodal understanding tasks under LLaVA-v1.5 setting. (2): We develop a unified MLLM architecture that demonstrates superior performance across both multimodal understanding and generation tasks, validating its versatility. (3): Our approach further pushes the performance boundary of unified discrete MLLMs, enabling improved scalability and representation capacity within next-token prediction frameworks.
2 Related Work
2.1 Image Tokenizer
Tokenization for Generation. Image tokenizers are crucial for autoregressive image generation[47; 40; 46]. VQVAE[47] learns a discrete representation using a learnable codebook in auto-encoder architectures. VQGAN[61] further improves better perceptual quality by using adversarial training[14]. advanced the existing architecture by integrating perceptual loss and discriminator loss, alongside adversarial loss, to enhance reconstruction quality. This approach yields more precise and detailed image representations, significantly improving upon previous methodologies in image generation and processing. Subsequently, ViT-VQGAN[62] and Efficient-VQGAN[5] advance the framework with the transformer design. In recent literature, researchers are turning to efficient codebook structures[41; 67; 62; 2] and better quantization methods[65; 63] to improve generation performance and compression rates. IBQ[41] propose the Index Backpropagation Quantization codebook update method, achieving stable training of large-scale codebooks. Although these methods efficiently retain low-level texture information, they frequently fail to capture high-level semantic information, which limits their application to multimodal understanding tasks.
Tokenization for Understanding. In multimodal large language models (MLLMs)[25; 24; 38; 30; 1; 9], researchers leverage CLIP[38] and BLIP[24] to extract visual characteristics that align with the language during its pre-training phase. Building upon, many works[30; 1; 9] have been collected and trained on high-quality datasets to achieve remarkable performance. LLaVA[30] utilizes vision encoder to align the vision inputs before LLMs. QwenVL[1] and InterVL[9] achieve better results through increased resolution, higher-quality datasets, etc. However, these text-aligned image encoders tend to focus only on semantic information and ignore texture information, which is important for the generation task.
2.2 Unified Image Tokenizer
Numerous efforts have emerged to develop unified visual generation and understanding within one MLLM[48; 50; 56; 11; 13; 44; 45; 55]. There are two main lines to bridge the gap. Many workers[11; 13; 44] combine diffusion models with LLMs for image generation. DREAMLLM[11] presents a unified framework that not only provides multimodal understanding but also creates multimodal content via diffusion models. Emu2[44] trains a unified generative model using a diffusion-based decoder. These approaches inevitably increase model complexity and are not simple enough. Other workers[45; 55; 56; 50] adopt VQVAE-based encoders to convert images into discrete tokens. Chameleon[45] and EMU3[50] directly use VQGAN[61], which is optimized by pixel reconstruction as the image tokenizer, while this method increases resource consumption during the pre-training stage and degrades multimodal understanding performance. VILA-U[55] introduce a unified image tokenizer that incorporates a text-aligned branch within the VQGAN training paradigm. However, due to the gap between semantic feature and texture feature, the joint optimization approach may lead to suboptimal solutions In contrast, our proposed SemHiTok can add the ability of extracting texture features without changing the discrete semantic features, and avoids the challenges brought by joint optimization.
@@ 3 Method
Figure 3: (a) SemHiTok is structurally composed of two branches: semantic branch and pixel branch. The semantic branch is trained following the VQKD[51], where the semantic codebook is learned through semantic loss. In the pixel branch, we propose semantic-guided hierarchical codebook(SGCH) composed of multiple pixel sub-codebooks, in which each pixel sub-codebook is in a one-to-one correspondence with a semantic code. The selection of pixel sub-codebook is indexed by the semantic code from semantic quantization. (b) Each semantic code is allocated to corresponding pixel sub-codebook, and concatenate their features along the dimension. (c) An illustration of the unified MLLM framework supporting both multimodal understanding and image generation.
The main objective of SemHiTok is to establish a simple and unified image tokenizer for multimodal understanding and generation. In this model, the image is transformed into discrete tokens that contain high-level semantic information and fine-grained texture information. We start with a semantic codebook training recipe that reconstructs semantic features extracted from language image pre-training model[66; 51], and point out semantic codebnook’s poor texture feature representation in section 3.1. In section 3.2, we conduct a preliminary discussion and observation. Building on this observation, we introduce Semantic-Guided Hierarchical Codebook(SGHC), incorporating texture information while perfectly inheriting the semantic information of semantic codebook, to enable pixel reconstruction enablement. Finally, we introduce the application of SemHiTok on unified MLLM in section 3.3. The overview framework is shown in Fig.3.
@@ 3.1 Semantic Codebook Training
For multimodal understanding, use text-aligned visual encoder[66; 37; 43; 49] as image tokenizer can accelerate convergence and improve performance. However, these text-aligned visual encoder typically output continuous semantic features. In this work, to achieve a unified visual tokenizer, we first train a semantic codebook to quantize the continuous semantic feature following VQKD[51].
@@
Where $\mathcal{E}_{\text{sem}}$ is a frozen text-aligned image encoder, e.g., CLIP[37] or SigLIP[66]. Then $Z_{sem}$ are transformed into discrete feature space $\mathcal{C}_{sem}=\{c_{1},c_{2},...,c_{K}\}\in\mathbb{R}^{K\times d_{sem}}$ through quantization function $\mathcal{Q}_{sem}(*)$. The quantization process $\mathcal{Q}_{sem}(*)$ is as follows:
@@
For $\mathcal{C}_{sem}$, we adopt EMA[17] VQ as the semantic codebook. Unlike traditional quantization methods, the EMA VQ is not updated via gradient descent, but instead through an Exponential Moving Average (EMA) algorithm:
@@
where $\mathbf{c}_{k}^{t}$ denotes the $k$-th codebook vector at update step $t$, $m$ is the momentum term, and the update is based on the average of all input vectors $\mathbf{z}_{i}$ assigned to code $\mathbf{k}$ in the current batch. We conduct multimodal understanding experiments using discrete tokens from $\mathcal{C}_{sem}$ in LLaVA-v1.5 setting as shown in Tab.4. The experiments show that our semantic codebook $\mathcal{C}_{sem}$ is slightly inferior to continuous features but still viable. However, we further conduct an experiment that reconstruct the original pixel from the quantized semantic features extracted by $\mathcal{C}_{sem}$. The reconstructed images exhibited noticeable blurriness and a significant loss of high-frequency details, as shown in Fig. 5. This indicates that the semantic codebook lacks pixel information.
Figure 4: Visualization of reconstructed by quantized semantic features. Semantic codebook tends to produce inaccuracies in the reconstruction of image textures and color information.
@@ 3.2 Pixel Reconstruction Enablement
Discussion. In section 3.1, we demonstrate that semantic code lacks the ability to model pixel information. In order to pixel reconstruction enablement and avoid a reduction of understand ability, a straightforward approach is to add an extra VQGAN[61] model. Semantic codebook extracts discrete semantic tokens for multimodal understanding, and VQGAN extracts discrete texture tokens for generation. The two token sets are concatenated—either dimensionally or sequentially, and passed to the LLM. However, the resulting token inflation or oversized codebook introduces significant computational burden, limiting its feasibility for MLLMs.
Furthermore, we present the visualization results of the semantic code, as shown in Fig.5. It can be observed that image patches corresponding to the same code exhibit similar pixel features. For example, the code $\mathit{v}_{14312}$ is more likely to be assigned to the rooster comb element in the image. At the same time, the image patches corresponding to these combs exhibit similar pixel features, such as color, patterns, and shapes. Based on this observation, we propose Semantic-Guided Hierarchical Codebook to model the pixel feature space corresponding to each semantic code using sub-codebook.
Semantic-Guided Hierarchical Codebook (SGHC). The SGHC consists of pretrained semantic codebook and several sub-codebooks, where each sub-codebook corresponds to a semantic code of semantic codebook, as shown in Fig. 3 (a). Specially, given the pre-trained semantic codebook $\mathcal{C}_{sem}=\{c_{1},c_{2},...,c_{K}\}\in\mathbb{R}^{K\times d_{sem}}$, the pixel codebook $\mathcal{C}_{pix}=\{\mathcal{C}_{pix}^{1},\mathcal{C}_{pix}^{2},\dots,\mathcal{C}_{pix}^{K}\}\in\mathbb{R}^{K\times m\times d_{pix}}$, where $\mathcal{C}_{pix}^{k}\in\mathbb{R}^{m\times d_{pix}}$ is $k_{th}$ semantic code’s pixel sub-codebook, $m$ is sub-codebook size. At first, the semantic codebook quantizes $X$ to discrete semantic token $Z_{sem}$ and semantic codebook index $I_{sem}$. In parallel, pixel encoder $\mathcal{E}_{pix}$ extract continuous pixel features $Z_{pix}=\mathcal{E}_{pix}(X)$. For the quantization process of the pixel codebook, the corresponding pixel sub-codebook is selected based on the quantization result of the semantic codebook. For instance, given image patch $i$, its semantic quantization codebook index $k$ and continuous pixel feature $Z_{pix}^{i}$, SGHC selects pixel sub-codebook $\mathcal{C}_{pix}^{k}$ to quantize $Z_{pix}^{i}$. The process is as follows:
@@
where $j$ is selected sub-codebook internal index. Finally, the semantic quantized tokens and pixel quantized tokens are concatenated to $Z_{q}=concate_{dim}(Z_{q_{sem}},Z_{q_{pix}})$ as the input of pixel decoder $\mathcal{D}_{pix}$ to reconstruct raw pixel image:
@@
where $\hat{X}$ is reconstructed pixel image. The $\mathcal{E}_{pix}$, $C_{pix}$ and $\mathcal{D}_{pix}$ are end-to-end trainable by minimizing reconstruction loss $L_{img}=\ell_{1}(\hat{X},X)$, codebook loss $Lc$, perceptual loss $L_{per}$ and represents adversarial loss $L_{gan}$[61]. The reconstruction loss is formulated as:
@@
Our SGHC can be regarded as the refinement of semantic discrete space to enable pixel reconstruction. We place a specific emphases on two key advantage of SGHC: (1) Non-Conflicting Extension: Our method leverages a pre-trained semantic codebook as foundation, with pixel reconstruction losses exclusively employed to optimize pixel branch modules during the PRE. This strategic approach effectively circumvents the suboptimal solutions that arise from joint optimization processes. Furthermore, SGHC’s final output is generated by concatenating semantic-quantized features with texture-quantized features, preserving the full expressive capacity of the original semantic features while integrating complementary texture information through this unified feature fusion paradigm. Subsequent tasks such as reconstruction, multimodal understanding, and generation are all share the same discrete token representation; (2) Efficient Downstream Applications: SGHC effectively avoids two critical predicaments: token quantity inflation and codebook overexpansion. As defined before, semantic codebook size is $K$ and each pixel sub-codebook size is $m$. Due to dimensional concatenation, the complete codebook flattens to $K\times m$, where $m$ is much more smaller than $K$. In our experimental default settings, we extend complete codebook to a size comparable to existing LLM’s vocabulary size, e.g., the size of Qwen2 vocabulary [60; 49] is $150$k. In addition, unlike VILA-U[54], our approach supports unified text and visual head, allowing cross-modal modeling. Leveraging these strengths, our framework demonstrates robust performance across multimodal downstream tasks, validating the effectiveness of the proposed architectural design.
Table 2: Comparison of reconstruction quality on the ImageNet-50k validation set. : quantizer uses residual quantization (RQ), where the total Code Num are multiplied by RQ depth. : quantizer uses multiple codebooks and product quantization.
@@
LlamaGen[42] 256 256 16,384 2.19
RQVAE[19] 256 256 16,384 3.20
RQVAE[19] 256 1024 16,384 1.30
VQGAN-LC[68] 256 256 16,384 3.01
VQGAN-LC[68] 256 256 100,000 2.62
IBQ[41] 256 256 16,384 1.37
IBQ[41] 256 256 262,144 1.00
FQGAN[2] 256 256 $16384\times 2$ 0.94
@@
VILA-U[54] 256 1024 16,384 1.80
SDE[57] 256 256 32,768 2.26
TokenFlow[36] 256 680 32,768 1.37
TokLIP[28] 256 256 16,384 2.19
UniTok[34] 256 256 $16384\times 4$ 0.39
SemHiTok(ours) 256 256 196,608 1.16
SemHiTok(ours) 384 729 196,608 0.66
Figure 6: Visualized reconstruction results from the ablation of key modules. PRE brings about a significant improvement in reconstruction quality compared to using only semantic discrete features. Moreover, the Enhance Decoder(ED) further improves reconstruction on hard samples, such as human faces and fine-grained textures.
3.3 Unified MLLM
The framework diagram for unified MLLM is shown in Fig.3(c). We use SemHiTok to develop a unified multimodal model, which models discrete vision and text token sequences with a universal next-token prediction loss. Particularly, in the image processing, SemHiTok is utilized to discretize images into token sequences. On the model side, we merely expand text vocabulary and adjust the head layer to incorporate visual token IDs. To enable a unified head layer, we flatten SGHC by merging all sub-codebooks into a single flat representation as shown in Fig.3(b). Specifically, for the $j_{th}$ semantic code in the $i_{th}$ pixel sub-codebook, the discrete code index in the completed codebook is $h=i\times m+j$, where $m$ is sub-codebook size. It is also worth noting that the vocabulary expansion is merely for implementation convenience. We still use the features extracted from SGHC as input and align with LLM through a lightweight adapter layer. To enable LLM to better handle features at two different levels, we introduce a Dual-MLP adapter layer, which projects semantic features and pixel features separately, and then concatenates them along the dimension before feeding them into the LLM. To enable classifier-free guidance[15], we randomly replace the text condition with probability of 0.1 to the un-condition text during training.

### v5

# arXiv:2503.06764v5[cs.CV] 28 Feb 2026

@@ Abstract
In this paper, we introduce SemHiTok, a unified image Tokenizer via Semantic-Guided Hierarchical codebook that provides consistent discrete representations for multimodal understanding and generation. Recently, unified image tokenizers have sparked exploration within the research community, which is designed to capture high-level semantic features for understanding and retaining low-level pixel features for generation. Previous works attempt to train a unified image tokenizer by combining loss for semantic distillation and pixel reconstruction. However, due to the differing levels of features prioritized by multimodal understanding and generation, joint training methods face significant challenges in achieving a good trade-off. SemHiTok addresses this challenge through a novel semantic-guided hierarchical codebook, which builds pixel sub-codebooks on a pretrained semantic codebook. This design decouples the semantic and pixel in terms of structure and training strategy, enabling the tokenizer to capture pixel features while retaining its ability to comprehend high-level semantic information. Our experiments demonstrate that SemHiTok achieves leading performance in image reconstruction and multimodal understanding under the LLaVA-v1.5 setting. Further, we develop a unified MLLM with SemHiTok, which exhibits superior performance across multimodal understanding and generation tasks. Extensive experiments confirm our analysis, showing that our unified image tokenizer architecture achieves a better trade-off.
@@ 1 Introduction
In recent years, autoregressive models have achieved great success in natural language processing and have been extended to the multimodal understanding domain, demonstrating immense potential. This triggers researchers’ interest in unified multimodal understanding and generation by employing a single autoregressive framework. To achieve a unified multimodal large model, the key challenge is designing a tokenizer suitable for both multimodal generation and understanding tasks.
@@
However, there is a vast gap in the visual information required for these two tasks. For instance, models from the CLIP (Radford et al., 2021a; Sun et al., 2023; Zhai et al., 2023) family, commonly used in multimodal understanding tasks, tend to lose visual pixel information. On the contrary, the VQGAN (Yu et al., 2021a; Zhu et al., 2024a) family models, often used in autoregressive generation tasks, lack the ability to extract semantic features for multimodal understanding tasks. This leads to poor performance when a single tokenizer is applied to a unified MLLM (Wu et al., 2024b; Jin et al., 2023; Li et al., 2024b). In light of the aforementioned issues, some recent work has attempted to incorporate a semantic learning branch into the original VQGAN training pipeline, aiming to obtain a unified tokenizer via joint optimization. VILA-U (Wu et al., 2024c) employs a straightforward combination of semantic alignment loss and pixel reconstruction loss, which allows the model to capture both low-level and high-level information. Nevertheless, the reliance on a hybrid structure and joint optimization often drives the tokenizer toward a suboptimal solution. Furthermore, some recent works (Xie et al., 2025; Qu et al., 2024) have built upon it with further improvements. As shown in Fig.1(a), while SDE (Xie et al., 2025) further decouples the encoders, the remaining hybrid codebook continues to impede the model’s optimization. TokenFlow (Qu et al., 2024) uses shared mapping to decouple the semantic branch and pixel branch while maintaining the consistency of the codebook index, but joint training still affects the final performance.
@@
A straightforward approach is to use CLIP and VQGAN to extract semantic and pixel information, respectively, and the concatenation of these two token sequences is then used as a unified representation. Janus (Wu et al., 2024a) introduces a dual-encoder method that separates encoders for understanding and generation tasks to address this conflict, but this increases the complexity of handling mixed tasks and does not fundamentally resolve the feature conflict challenge. However, this leads to a doubling of the token sequence count or multiplicative expansion in vocabulary size, depending on whether the concatenation is applied along the length or dimension. These limitations underscore a fundamental challenge in the field: How to balance semantic-level and pixel-level information effectively, without compromising the ease of integration into MLLM frameworks?
Figure 1: Illustration of other tokenizers and SemHiTok. HyStruct: Using a single model to extract information at different levels; DStruct: Using different models to extract information at various levels; JTrain: Using a joint optimization training strategy; DTrain: Adopt a phased optimization training strategy.
@@
To address this challenge, we propose SemHiTok, a unified image tokenizer that provides consistent feature representations for multimodal understanding and generation tasks through a unique hierarchical codebook design. Inspired by the observation that image patches with the same semantic code tend to have similar pixel features, we introduce a novelty hierarchical codebook which uses a sub-codebook to model the pixel-level space associated with each semantic code, named Semantic-Guided Hierarchical Codebook(SGHC). Unlike existing approaches, SemHiTok supports a stage-wise training paradigm where each stage exclusively optimizes specific hierarchy level features, allowing us to achieve a better trade-off between semantic and pixel feature extraction. In addition, SemHiTok can be seamlessly integrated into existing MLLMs following the next-token paradigm through a simple codebook flattening operation.
@@
Our contributions can be summarized as follows: (1): A novel unified tokenizer that achieves a trade-off between semantic and pixel information, demonstrating outstanding performance in both image reconstruction and multimodal understanding tasks. (2): We develop a unified MLLM architecture that demonstrates superior performance across both multimodal understanding and generation tasks, validating its versatility. (3): Our approach further pushes the performance boundary of unified discrete MLLMs, enabling improved scalability and representation capacity within next-token prediction frameworks.
@@ 2 Method
Figure 2: (a) SemHiTok is structurally composed of two branches: semantic branch and pixel branch. The semantic branch is trained following the VQKD (Wei et al., 2022), where the semantic codebook is learned through semantic loss. We propose a semantic-guided hierarchical codebook(SGCH) composed of multiple pixel sub-codebooks, in which each pixel sub-codebook is in a one-to-one correspondence with a semantic code. The selection of pixel sub-codebook is indexed by the semantic code from semantic quantization. To enable a unified discrete representation, we concatenate the quantized semantic and pixel features along the channel dimension and feed the result into the pixel decoder for reconstruction. (b) Each semantic code is allocated to the corresponding pixel sub-codebook, and their features are concatenated along the dimension. (c) An illustration of the unified MLLM framework.
The main objective of SemHiTok is to establish a simple and unified image tokenizer for multimodal understanding and generation. In this model, the image is transformed into discrete tokens that contain semantic information and pixel information. We begin with a semantic codebook training recipe that reconstructs semantic features extracted from a language image pre-training model (Zhai et al., 2023; Wei et al., 2022), and point out the semantic codebook’s poor texture feature representation in section 2.1. In section 2.2, we conduct a preliminary discussion and observation. Building on this observation, we introduce Semantic-Guided Hierarchical Codebook(SGHC), incorporating texture information while perfectly inheriting the semantic information of the semantic codebook, to enable pixel reconstruction enablement. Finally, we introduce the application of SemHiTok on unified MLLM in section 2.3. The overview framework is shown in Fig.2.
@@ 2.1 Semantic Codebook Training
For multimodal understanding, using a text-aligned visual encoder (Zhai et al., 2023; Radford et al., 2021a; Sun et al., 2023; Wang et al., 2024b) as an image tokenizer can accelerate convergence and improve performance. However, these text-aligned visual encoders typically output continuous semantic features. In this work, to achieve a unified visual tokenizer, we first train a semantic codebook to quantize the continuous semantic feature following VQKD (Wei et al., 2022).
@@
Where $\mathcal{E}_{\text{sem}}$ is a frozen text-aligned image encoder, e.g., CLIP (Radford et al., 2021a) or SigLIP (Zhai et al., 2023). Then $Z_{sem}$ are transformed into discrete feature space $\mathcal{C}_{sem}=\{c_{1},c_{2},...,c_{K}\}\in\mathbb{R}^{K\times d_{sem}}$ through quantization function $\mathcal{Q}_{sem}(*)$. The quantization process $\mathcal{Q}_{sem}(*)$ is as follows:
@@
For $\mathcal{C}_{sem}$, we adopt EMA (Hunter, 1986) VQ as the semantic codebook. Unlike traditional quantization methods, the EMA VQ is not updated via gradient descent, but instead through an Exponential Moving Average (EMA) algorithm:
@@
where $\mathbf{c}_{k}^{t}$ denotes the $k$-th codebook vector at update step $t$, $m$ is the momentum term, and the update is based on the average of all input vectors $\mathbf{z}_{i}$ assigned to code $\mathbf{k}$ in the current batch. However, we further conduct an experiment that reconstructs the original pixel from the quantized semantic features extracted by $\mathcal{C}_{sem}$. The reconstructed images exhibited noticeable blurriness and a significant loss of high-frequency details, as shown in Fig. 4. This indicates that the semantic codebook lacks pixel information.
Figure 3: Visualization of reconstruction by quantized semantic features. Semantic codebook tends to produce inaccuracies in the reconstruction of image textures and color information.
@@ 2.2 Pixel Reconstruction Enablement
Discussion. In section 2.1, we demonstrate that semantic code lacks the ability to model pixel information. In order to pixel reconstruction enablement and avoid a reduction of understand ability, a straightforward approach is to add an extra VQGAN (Yu et al., 2021a) model. Semantic codebook extracts discrete semantic tokens for multimodal understanding, and VQGAN extracts discrete texture tokens for generation. The two token sets are concatenated—either dimensionally or sequentially, and passed to the LLM. However, the resulting token inflation or oversized codebook introduces a significant computational burden, limiting its feasibility for MLLMs.
Furthermore, we present the visualization results of the semantic code, as shown in Fig.4. It can be observed that image patches corresponding to the same code exhibit similar pixel features. For example, the code $\mathit{v}_{14312}$ is more likely to be assigned to the rooster comb element in the image. At the same time, the image patches corresponding to these combs exhibit similar pixel features, such as color, patterns, and shapes. Based on this observation, we propose Semantic-Guided Hierarchical Codebook to model the pixel feature space corresponding to each semantic code using a sub-codebook.
Semantic-Guided Hierarchical Codebook (SGHC). The SGHC consists of a pretrained semantic codebook and several sub-codebooks, where each sub-codebook corresponds to a semantic code of the semantic codebook, as shown in figure 2 (a). Specially, given the pre-trained semantic codebook $\mathcal{C}_{sem}=\{c_{1},c_{2},...,c_{K}\}\in\mathbb{R}^{K\times d_{sem}}$, the pixel codebook $\mathcal{C}_{pix}=\{\mathcal{C}_{pix}^{1},\mathcal{C}_{pix}^{2},\dots,\mathcal{C}_{pix}^{K}\}\in\mathbb{R}^{K\times m\times d_{pix}}$, where $\mathcal{C}_{pix}^{k}\in\mathbb{R}^{m\times d_{pix}}$ is $k_{th}$ semantic code’s pixel sub-codebook, $m$ is sub-codebook size. At first, the semantic codebook quantizes $X$ to a discrete semantic token $Z_{sem}$ and a semantic codebook index $I_{sem}$. In parallel, pixel encoder $\mathcal{E}_{pix}$ extract continuous pixel features $Z_{pix}=\mathcal{E}_{pix}(X)$. For the quantization process of the pixel codebook, the corresponding pixel sub-codebook is selected based on the quantization result of the semantic codebook. For instance, given image patch $i$, its semantic quantization codebook index $k$ and continuous pixel feature $Z_{pix}^{i}$, SGHC selects pixel sub-codebook $\mathcal{C}_{pix}^{k}$ to quantize $Z_{pix}^{i}$. The process is as follows:
@@
where $j$ is the selected sub-codebook internal index. Finally, the semantic quantized tokens and pixel quantized tokens are concatenated to $Z_{q}=concate_{dim}(Z_{q_{sem}},Z_{q_{pix}})$ as the input of pixel decoder $\mathcal{D}_{pix}$ to reconstruct the raw pixel image:
@@
where $\hat{X}$ is reconstructed pixel image. The $\mathcal{E}_{pix}$, $C_{pix}$ and $\mathcal{D}_{pix}$ are end-to-end trainable by minimizing reconstruction loss $L_{img}=\ell_{1}(\hat{X},X)$, codebook loss $Lc$, perceptual loss $L_{per}$ and represents adversarial loss $L_{gan}$ (Yu et al., 2021a). The reconstruction loss is formulated as:
@@
Our SGHC can be regarded as the refinement of a semantic discrete space to enable pixel reconstruction. We place a specific emphasis on two key advantages of SGHC: (1) Non-Conflicting Extension: Our method leverages a pre-trained semantic codebook as a foundation, with pixel reconstruction losses exclusively employed to optimize pixel branch modules during the PRE. This strategic approach effectively circumvents the suboptimal solutions that arise from joint optimization processes. Furthermore, SGHC’s final output is generated by concatenating semantic-quantized features with pixel-quantized features, preserving the full expressive capacity of the original semantic features while integrating complementary texture information through this unified feature fusion paradigm. Subsequent tasks, such as reconstruction, multimodal understanding, and generation, all share the same discrete token representation; (2) Efficient Downstream Applications: SGHC effectively avoids two critical predicaments: token quantity inflation and codebook overexpansion. As defined before, the semantic codebook size is $K$, and each pixel sub-codebook size is $m$. Due to dimensional concatenation, the complete codebook flattens to $K\times m$, where $m$ is much smaller than $K$. In our experimental default settings, we extend the complete codebook to a size comparable to existing LLMs’ text vocabulary size, e.g., the size of Qwen2 vocabulary is 150150k.
Table 1: Comparison of reconstruction quality on the ImageNet-50k validation set. : quantizer uses residual quantization (RQ), where the total Code Shape is multiplied by RQ depth. : quantizer uses multiple codebooks and product quantization.
@@
LlamaGen (Sun et al., 2024a) 256 $16\times 16$ 16,384 2.19
RQVAE (Lee et al., 2022) 256 $16\times 16\times 4$ 16,384 3.20
VQGAN-LC (Zhu et al., 2024a) 256 $16\times 16$ 100,000 2.62
IBQ (Shi et al., 2024) 256 $16\times 16$ 16,384 1.37
IBQ (Shi et al., 2024) 256 $16\times 16$ 262,144 1.00
FQGAN (Bai et al., 2024) 256 $16\times 16\times 2$ $16,384\times 2$ 0.94
@@
VILA-U (Wu et al., 2024c) 256 16x16x4 16,384 1.80
SDE(MUSE-VL) (Xie et al., 2025) 256 $16\times 16$ 32,768 2.26
TokenFlow (Qu et al., 2024) 256 680 32,768 1.37
TokLIP (Lin et al., 2025) 256 $16\times 16$ 16,384 2.19
QLIP-B (Lin et al., 2025) 256 $16\times 16$ $2^{28}$ 3.21
UniTok (Ma et al., 2025) 256 $16\times 16$ $16,384\times 4$ 0.39
SemHiTok(ours) 256 $16\times 16$ 196,608 1.16
SemHiTok(ours) 384 $27\times 27$ 196,608 0.66
Figure 5: Visualized reconstruction results from the ablation of key modules. PRE brings about a significant improvement in reconstruction quality. Moreover, the Enhance Decoder(ED) further improves reconstruction on hard samples.
2.3 Unified MLLM Equipped with SemHiTok
The framework diagram for unified MLLM is shown in Fig.2(c). We use SemHiTok to develop a unified multimodal model, which models discrete vision and text token sequences with a universal next-token prediction loss. Particularly, in image processing, SemHiTok is utilized to discretize images into token sequences. On the model side, we merely expand the text vocabulary and adjust the head layer to incorporate visual token IDs. To enable a unified head layer, we flatten SGHC by merging all sub-codebooks into a single flat representation as shown in Fig.2(b). Specifically, for the $j_{th}$ semantic code in the $i_{th}$ pixel sub-codebook, the discrete code index in the completed codebook is $h=i\times m+j$, where $m$ is the sub-codebook size. It is also worth noting that the vocabulary expansion is merely for implementation convenience. We still use the features extracted from SGHC as input and align with LLM through a lightweight adapter layer. After training is completed, we replace the visual component in the embedding layer in order to achieve consistency between training and inference. To enable LLM to better handle features at two different levels, we introduce a Dual-MLP adapter layer, which projects semantic features and pixel features separately, and then concatenates them along the dimension before feeding them into the LLM. To enable classifier-free guidance (Ho and Salimans, 2022), we randomly replace the text condition with a probability of 0.1 to the unconditioned text during training. More deployment details are provided in the Supplementary Material 7.2.